Hello '
This is a new question. Guy wire. How long reaching from the top of a 9 feet pole to a point on the ground 6 feet from the pole. What is its exact lenght in feet?

Answer by nyc_function(2741)

Again, use the Pythagorean Theorem.
Let g = guy wire length
9^2 + 6^2 = g^2
81 + 36 = g^2
117 = g^2
sqrt[117] = g
Did you follow?
NOTE: sqrt = square root for short.